OpenStudy (anonymous):

If the slope of the tangent to the curve x^2-2xy+4=0 at a point on it is -3/2, then find tangent and normal at that point (Applications of Differentiation) Can anyone help me with this?

OpenStudy (shubhamsrg):

Whats your attempt/progress ?

OpenStudy (anonymous):

I found out the Y derivative of the curve and equated it to -3/2 and I got another equation in x and y terms... I don't know what to do after that....

OpenStudy (shubhamsrg):

One eqn you'll get after differentiating, 1 you already have. You know the slope. You just need a point, solve both eqns to get the req point.

OpenStudy (anonymous):

but one eqn is linear and the other quadratic hmm shou

OpenStudy (anonymous):

should i factorize the given curve eqn

OpenStudy (anonymous):

oopa no that's not possible

OpenStudy (shubhamsrg):

2x - 2y - 2xdy/dx = 0 x-y -3x/2 = 0 y = -x/2 IS that what you get ?

OpenStudy (shubhamsrg):

oh wait, correction : it should be 5x=2y

OpenStudy (anonymous):

I AM SOO SORRY THE Question is this...... \[x ^{2}-2xy+4y=0\]

OpenStudy (shubhamsrg):

hmm, so what eqns you get ?

OpenStudy (anonymous):

After diff and euating it to -3/2.... this is what I got.... \[5x-2y=6\]

OpenStudy (shubhamsrg):

okay. seems right.

OpenStudy (shubhamsrg):

original eqn is x^2 - 2xy + 4y =0 its easier if you substitute for y from 1st eqn.

OpenStudy (shubhamsrg):

In case you are following ?

OpenStudy (anonymous):

I tried doing that but after substituting for Y the equation we get cannot be factorized directly... x^2+x-3=0

OpenStudy (shubhamsrg):

I am getting a solution, you have probably done mistake in substituing, tey again.

OpenStudy (shubhamsrg):

substituting* try**

OpenStudy (shubhamsrg):

problems ?

OpenStudy (anonymous):

hmm no am getting the same equation after substituting...

OpenStudy (shubhamsrg):

5x-2y =6 y = (5x-6)/2

OpenStudy (anonymous):

yes i substituted this in the original equation

OpenStudy (shubhamsrg):

x^2 - 2xy + 4y =0 x^2 - 2x(5x-6)/2 + 4(5x-6)/2 = 0 x^2 - x(5x-6) + 2(5x-6) = 0

OpenStudy (anonymous):

it should be 2x^2 in the last line

OpenStudy (shubhamsrg):

eh ?

OpenStudy (anonymous):

yea should take 2 as LCM and then send it to the other side, you forgot to multiply 2 to the first term x^2

OpenStudy (shubhamsrg):

I did not forget ,I cancelled out 2, one from "2"xy , other from "4"y

OpenStudy (shubhamsrg):

SO final eqn which you should be getting is -4x^2 + 16x - 12 = 0 or x^2 - 4x + 3 =0

OpenStudy (shubhamsrg):

And am sorry for my late replies, am generally not slow, just busy somewhere else also.

OpenStudy (anonymous):

ohh yeaaa right right :) thank you so much :) I got the points thanks you.... !!

OpenStudy (shubhamsrg):

Glad to help ^_^